OpenStudy (anonymous):

Find the greatest common factor of the following monomials: gh^5 10g^5h^4

OpenStudy (welshfella):

the GCF is the greatest term which will divide into the 2 mononomials

OpenStudy (welshfella):

take them letter by letter GCF of g and g^5 = ?

OpenStudy (welshfella):

GCF of h^4 and h^5 = ?

OpenStudy (anonymous):

g and h^4

OpenStudy (welshfella):

yes thats right

OpenStudy (welshfella):

so whats the GCF?

OpenStudy (anonymous):

um 1?

OpenStudy (welshfella):

no you just combine the 2 you just mentioned

OpenStudy (welshfella):

GCF = gh^4

OpenStudy (anonymous):

ahh okay! thank you

OpenStudy (welshfella):

yw